OpenStudy (anonymous):

On a circle of radius 6 feet, what angle would subtend an arc of length 5 feet? (Answer in Degrees)

Directrix (directrix):

The length of the subtended arc is to the circumference of the circle as the central angle that subtends that arc is to 360 degree, the analog in degrees to the circumference of the circle in feet. 5 is to 2 pi r as A is to 360 degrees where A is the measure of the angle 5/[2 pi (6)] = A / 360 5/ ( 12 pi) = A / 360 Cross multiply and solve for A. Let us know what you get and we can compare answers. Okay? @deerubio